OpenStudy (anonymous):

A carton of 12 rechargable batteries contains one that is defective. In how many ways can the inspector choose 3 of the batteries and (a) get the one that is defective, (b)not get the one that is defective?

OpenStudy (anonymous):

Can't seem to figure this one out. They get 55 for the first one.

OpenStudy (anonymous):

i guess for the first one .. that now we left with 11 batteries of which we have to take 2 of them .. so \[\left(\begin{matrix}11 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

oh okay, so \[\left(\begin{matrix}11 \\ 2\end{matrix}\right)=\frac{10*10}{2*1}\\]

OpenStudy (anonymous):

no \[\left(\begin{matrix}11 \\ 2\end{matrix}\right) = 11! / (2! * 9!)\]

OpenStudy (anonymous):

Oh so its \[\left(\begin{matrix}n \\ r\end{matrix}\right)=\frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay so I think I've figured out how they got the second one: \[\left(\begin{matrix}11 \\ 3\end{matrix}\right)=\frac{11!}{3!}=\frac{11*10*9}{3*2*1}=165\]

OpenStudy (anonymous):

So the restriction in this example,r, is 3 because I have three possibilities (by choosing three batteries at a time) of not getting the defective battery, where as in the previous scenario I had two possibilities . Is this correct?

OpenStudy (anonymous):

yes it is :)